How many milligrams of silver bromide would dissolve in 1.0 liter of water?
Answer #1

Ans:

Solubility of silver bromide (AgBr) in water at 293K = 0.14 mg/L

Meaning, 0.14 mg of AgBr is soluble in 1 L of water at 20oC.

OR

Solubility product (Ksp) of AgBr in water at 293 K = 5.4 × 10−13 (data collected)

Let ‘S M’ be the solubility of AgBr in water.

Then, Ksp = S2 for AgBr

            S = (Ksp)½ = (5.4 x10-13)½ = 7.348 x 10-7 M = 7.348 x 10-7 mol/L

Meaning, 7.348 x 10-7 mol of AgBr is soluble in 1 L of water.

Molar mass of AgBr = 187.77 g/mol

7.348 x 10-7 mol means = 7.348 x 10-7 mol x 187.77 g/mol = 0.000138 g = 0.14 mg